### v1

# ImpliRet: Benchmarking the Implicit Fact Retrieval Challenge

Abstract
Retrieval systems are central to many NLP pipelines, but often rely on surface-level cues such as keyword overlap and lexical semantic similarity. To evaluate retrieval beyond these shallow signals, recent benchmarks introduce reasoning-heavy queries; however, they primarily shift the burden to query-side processing techniques – like prompting or multi-hop retrieval – that can help resolve complexity. In contrast, we present ImpliRet, a benchmark that shifts the reasoning challenge to document-side processing: The queries are simple, but relevance depends on facts stated implicitly in documents through temporal (e.g., resolving “two days ago”), arithmetic, and world knowledge relationships. We evaluate a range of sparse and dense retrievers, all of which struggle in this setting: the best nDCG@10 is only 15.07%. We also test whether long-context models can overcome this limitation. But even with a short context of only ten documents, including the positive document, GPT-4.1 scores only 35.06%, showing that document-side reasoning remains a challenge. Our codes are available at github.com/ZeinabTaghavi/IMPLIRET.††* Equal Contribution.
ImpliRet: Benchmarking the Implicit Fact Retrieval Challenge
Zeinab Sadat Taghavi*, Ali Modarressi*, Yunpu Ma, Hinrich Schütze Center for Information and Language Processing, LMU Munich Munich Center for Machine Learning (MCML) Correspondence: zeinabtaghavi@cis.lmu.de
1 Introduction
Retrieval systems play a pivotal role in many NLP applications, enabling models to utilize relevant information from large corpora such as document collections, web pages, or conversational histories Lewis et al. (2020); Gao et al. (2023). Relevance in retrieval can be established through a range of connections, from explicit lexical or semantic similarity to more implicit, context-dependent associations. However, widely used retrieval systems are highly reliant on surface-level cues such as exact matches, repetition, or where a fact appears in the text Ram et al. (2023); Coelho et al. (2024); Fayyaz et al. (2025). Additionally, many popular benchmarks (e.g., BEIR (Thakur et al., 2021)) do not surface these issues as their queries have lexical overlap with relevant documents Shao et al. (2025). There are attempts to create reasoning-intensive datasets that push beyond lexical and surface-level matches. For instance, RAR-b Xiao et al. (2024) reframes multiple-choice reasoning tasks into retrieval problems, BIRCO Wang et al. (2024) collects multi-faceted questions across five domains, and BRIGHT Su et al. (2025) uses full StackExchange problem descriptions as queries against the pages they cite. Since the reasoning burden lies on the query side, techniques like query expansion, chain-of-retrieval inference, or agentic retrieval can help models handle complex prompts and outperform standard retrievers Wang et al. (2025); Song et al. (2025); Li et al. (2025).
In contrast, we present ImpliRet, a benchmark that shifts reasoning to document-side processing: the queries are simple, but relevance depends on facts stated implicitly within the documents, spanning arithmetic, temporal, and world knowledge relationships that require inference to uncover. Figure 1 gives an example: the correct document requires resolving a reference to a date that is implicit, i.e., not stated directly. An effective retrieval system must infer such implicit facts from the document content, ideally as part of the indexing process, in order to retrieve the correct result at query time. Yet current retrieval methods fail to capture the implicit signals needed for accurate retrieval. We evaluate sparse and dense approaches, including BM25 Robertson and Zaragoza (2009), ColBERT Santhanam et al. (2022), and Dragon+ Lin et al. (2023), and observe consistently poor performance: the best nDCG@10 is only 15.07% across our benchmark. To test whether long-context capabilities could mitigate the problem, we evaluate models in a setting where the positive document is included among several distractors. While GPT-4.1 answers correctly when given only the positive document, its performance drops sharply even with just ten documents in-context, achieving a ROUGE-1 recall of 35.06%. Our dataset ImpliRet introduces a new setting that requires document-side reasoning for retrieval rather than query-side reasoning. ImpliRet presents challenges for both retrieval and long-context processing, highlighting the need for models that can reason over implicit information embedded in large corpora.
2 ImpliRet
In ImpliRet, we construct examples whose relevance depends on information that is implicitly stated in the document, i.e., it can only be discovered through reasoning, not by surface-level overlap. ImpliRet covers three reasoning categories: World Knowledge, Arithmetic, and Temporal.
We compile a collection of implicit-tuple sets. Within each set, a tuple links an implicit surface form that appears in a document to the explicit form that will appear in the query; see Fig. 1, e.g. (“2024-10-13 … seven days ago”, “October 06, 2024”).
For every reasoning category, we create such tuple sets. Each set () contains unique tuples (). Tuples in the tuple sets are unique but not guaranteed to be unique throughout the collection of tuple sets. Hence, before document generation, we inject distinct auxiliary lexical entities (e.g. named entities, speaker names) into each tuple so that the documents generated from remain distinguishable from those of when (see Appendix A.4).
From each tuple in the tuple set, we generate a document, yielding a pool of documents with . The document derived from is the only positive for the query constructed from , whereas all other documents in the global collection – including those from tuples in the same set and every document from any other set – are treated as negatives.
For each reasoning category, we generate two collections of tuple sets, one realized in the uni-speaker style and the other in the multi-speaker style, keeping their respective document pools separate to foster surface diversity. Thus, every query has exactly one positive document, while every other document in the global collection serves as a semantically irrelevant negative. In the remainder of this section, we detail the construction of the implicit-tuple sets and our procedure for generating documents and queries.
2.1 Generating Tuple Sets
Arithmetic.
An arithmetic relation requires simple numerical reasoning. For instance, the query “Which bag costs $1,600?” can be answered by “The Prada bag costs $2,000, the Gucci bag is 20% cheaper,” since $2,000 × 0.8 = $1600. Here, the model must identify the reference price, interpret the relative statement (“20% cheaper”), and perform the corresponding computation to infer the answer. Therefore, each tuple in the implicit tuple set takes the form , where is the base price, is the relative multiplier, indicates the direction of the change, and is the queried price (e.g., ). We apply constraints to ensure that queried prices are unique, realistic, and well-distributed across the tuple set. Tuples are generated using a sampling algorithm that selects base prices and checks constraint satisfaction, backtracking as needed until valid tuples are found (where is the target number of documents indicated as “Docs” in Table 1). Full constraint details and sampling logic are provided in Appendix 2.1.
World Knowledge.
A world knowledge relation connects a textual mention to an external fact. For instance, the query “Who was in the UK?” can be answered by “Lenna was at Big Ben,” based on the implicit fact that Big Ben is located in the UK. The model must identify the mentioned entity, retrieve the associated world fact, and use it to resolve the query. Each tuple is encoded as , e.g., . To build the tuple set, we collect landmark-country pairs that are unambiguous, globally unique, free of lexical cues revealing the country, and refer to specific rather than generic locations. Candidates are sourced from Wikidata Vrandečić and Krötzsch (2014) and filtered using LLMs, embedding similarity, and web search verification. Full filtering criteria, prompts, and implementation details are provided in Appendix A.2. Here, we again generate a set of tuples of each implicit tuple set.
Temporal.
A temporal relation involves reasoning over relative dates; we gave an example in Figure 1. The model must identify the reference date (2024-10-13), interpret the relative time expression (“seven days ago”), and compute the resulting absolute date (“2024-10-06”). Each example is represented as a tuple , where is the base date explicitly mentioned in the document, is a list of relative offsets (e.g. [“1 day after”, “2 days after”]), and is the list of resolved explicit dates (e.g., [“March 6th”, “March 7th”]). We generate such tuples under constraints that ensure date uniqueness, broad coverage across a fixed window, and realistic time offsets. Target date sequences are first sampled, then anchored to a base date to define relative expressions. The sampling algorithm verifies constraints and backtracks as needed until a valid set is found. Further details on constraints and sampling logic are provided in Appendix A.3.
Reasoning | Style | Docs | Tokens | |
Avg. | Total | |||
Arithmetic | Uni-speaker | 1500 | 553 | 830,098 |
Multi-speaker | 1500 | 142 | 213,750 | |
World- | Uni-speaker | 1500 | 471 | 707,047 |
Knowledge | Multi-speaker | 1500 | 168 | 253,337 |
Temporal | Uni-speaker | 1500 | 479 | 719,226 |
Multi-speaker | 1500 | 141 | 212,502 |
2.2 Document-Query Pairs
We generate a document-query pair from every fact tuple, realizing it in one of two styles: uni-speaker (multi-turn chat) or multi-speaker (forum thread).
Uni-speaker (multi-turn chat).
For each tuple, we create a short multi-turn dialogue. The same main conversant (e.g., “Alex”) appears in every dialogue within a tuple set and never appears in any other tuple sets. To keep the interactions natural, the second conversant’s name changes from one dialogue to the next. Depending on the reasoning category, the main conversant states which product they bought at a certain price (Arithmetic), mentions visiting a landmark (World Knowledge), or describes an activity that occurred on a specific date (Temporal). The query then targets the implicit fact contained in that statement: the product, person, or activity linked to the given price, country, or date.
Multi-speaker (forum thread, one post per user).
Each tuple set receives a single prompt that serves as the thread’s opening post. For that tuple set, we create a forum thread in which each post is authored by a different user, realizing one tuple, and all posts respond to the shared prompt. Thus, the thread mimics a discussion in which several users independently mention their purchase, visit, or scheduled activity, respectively. While the underlying actions mirror the uni-speaker setting, the query perspective shifts: instead of asking about an attribute of a known entity, it now asks which entity (product, person, or activity) satisfies a stated condition such as a price, location, or date.
Generation Pipeline.
In both styles, i.e., in each conversation and post, every message includes a timestamp and speaker name (see Figure 1). In both styles, each example is produced via a three-step pipeline: (1) Entity binding: We assign entities (e.g., names, items, activities) to each tuple to create a plausible scenario and define the query target; (2) Document generation: We prompt an LLM to generate a chat or forum passage that embeds the entity and the implicit part of the tuple, without stating the explicit fact; (3) Verification: a second model attempts to extract the original tuple; we retain only examples where the intended fact is fully recoverable. This pipeline is supported by auxiliary lexical resources, including random names, brand-item pairs, and activity lists, as well as per-reasoning category prompt templates. We use LLAMA 3.3-70B Meta (2024) to synthesize the documents for each tuple.111Details such as prompts and query templates are available in Appendix A.4. Table 1 presents ImpliRet statistics222The tokens are counted using GPT-2 tokenizer Radford et al. (2019)..
3 Experiments
We employ ImpliRet to probe whether state-of-the-art retrievers can perform document-side reasoning. Relevant documents are retrieved for each query among those documents that are in its corresponding (reasoning category and discourse style) group.
At test time, each query is compared to all its discourse style documents. Our evaluation covers a wide variety of retrieval methods: sparse lexical baseline BM25 Robertson and Zaragoza (2009); Lù (2024); dense encoders Contriever, Dragon+, and ReasonIR Izacard et al. (2021); Lin et al. (2023); Shao et al. (2025); late interaction model ColBERT v2 Santhanam et al. (2022); and knowledge graph augmented retriever HippoRAG 2 Gutiérrez et al. (2025). Effectiveness is reported as nDCG@ in the main text; MRR@ appears in Appendix B.
Retriever | Reasoning | Average | ||
W. Know. | Arithmetic | Temporal | ||
Sparse | ||||
BM25 | 14.10 | 11.06 | 11.22 | 12.13 |
Late-Interaction | ||||
ColBERT v2 | 16.04 | 14.93 | 12.20 | 14.39 |
Dense Encoders | ||||
Contriever | 16.15 | 13.84 | 12.59 | 14.19 |
Dragon+ | 17.15 | 14.61 | 12.53 | 14.76 |
ReasonIR | 19.53 | 10.74 | 14.93 | 15.07 |
Knowledge-Graph Augmented Indexer | ||||
HippoRAG 2 | 16.38 | 14.21 | 12.50 | 14.36 |
4 Results
The nDCG@10 results across all reasoning categories are presented in Table 2. The highest average score, 15.07 (achieved by ReasonIR, a recent 8B-parameter LLM), shows the difficulty retrieval models face when reasoning over implicit facts in documents. More efficient baselines such as Contriever, Dragon+, and BM25 perform substantially worse; notably, BM25 reaches just 12.13 due to its reliance on surface-level lexical overlap.
Performance varies across reasoning types: the World Knowledge category exhibits the largest performance spread (14.10 vs. 19.53), while it is narrowest for Arithmetic (10.74 vs. 14.61). Discourse style also plays a role: ReasonIR scores 20.58 on multi-speaker examples compared to 9.56 on uni-speaker ones, suggesting that stylistic structure affects retrieval difficulty.333Full results per category and style in Appendix B.
Experiment | Reasoning | Average | |||
W. Know. | Arithmetic | Temporal | |||
Llama 3.3 70B | 1 | 73.79 | 90.13 | 81.85 | 81.92 |
10 | 27.37 | 16.98 | 25.23 | 23.19 | |
All | 17.43 | 4.42 | 10.29 | 10.71 | |
GPT-4.1 | 1 | 93.24 | 92.12 | 84.90 | 88.05 |
10 | 62.21 | 23.86 | 15.59 | 35.06 | |
All | 53.91 | 9.28 | 6.93 | 22.90 |
RAG Performance with an Oracle Retriever on Reason-Sensitive Documents.
While retrieval quality clearly affects end-to-end performance, we ask whether an LLM with long-context capacity can still succeed once the relevant document is present. To test this, we use a retrieval-augmented generation (RAG) set-up with an oracle retriever, one that always includes the positive document in its top-. The model sees the question together with documents: one positive and hard negatives sampled from the same pool (among other samples), ensuring comparable style and topic. This configuration removes retrieval as a variable and isolates the LLM’s document-side reasoning ability.
We evaluate three settings: (positive only), (positive plus nine negatives), and a full-pool setting where all documents from the pool are provided as context. The model receives the query along with the sequence of documents and must generate an answer. We evaluate two reader models: LLAMA 3.3 70B and GPT-4.1.444Checkpoint: gpt-4.1-2025-04-14 In Table 3, we report the average ROUGE-1 recall555R-1 Rec. scores to measure the overlap between the generated output and the positive answer Lin (2004). When given only the positive document (), the two models achieve average ROUGE-1 Recall of 81.92 and 88.05. This suggests that the query itself is straightforward to answer once the relevant document is isolated. This also means that an LLM can solve the task if a high-performing retriever (which would retrieve the relevant document at rank 1) is available. However, as increases (even with the positive included), performance declines, showing that LLMs struggle to focus on the correct evidence amid structurally similar negatives. This supports prior findings on long-context limitations and highlights the need for retrieving a small, focused set of documents rather than increasing context size Kuratov et al. (2024); Modarressi et al. (2025).
5 Conclusion
We introduce ImpliRet, a benchmark for evaluating retrieval models when relevance depends on document-side reasoning on implicit facts. Unlike prior datasets that emphasize complex queries, ImpliRet shifts the reasoning burden to the documents. It covers three reasoning types – world knowledge, arithmetic, and temporal – and two discourse styles. Across sparse, dense, and KG-augmented retrievers, the best nDCG@10 is only 15.07. Even with GPT-4.1, given ten documents including the gold, performance peaks at just 35.06%. These results highlight the difficulty of retrieving implicit facts and the need for models that can reason beyond surface cues.
Limitations
While our benchmark is carefully designed to evaluate implicit document-side reasoning in retrieval systems, it has the following limitations:
Synthetic Dataset.
Documents and queries in ImpliRet are synthesized using LLMs and structured templates. This allows control over the facts and how they are implicitly expressed, while avoiding conflicts. It also enables easy regeneration if data contamination or memorization is suspected. As with any synthetic benchmark, the data may differ slightly from naturally occurring text in discourse structure or topic diversity. All examples are in English and follow conversational formats (uni-speaker chats and multi-speaker forum posts). Although the use of LLMs helps ensure fluency, it introduces the risk of subtle hallucinations or unintended cues, which we address through automatic verification during dataset construction.
Reasoning Types & Level.
In ImpliRet, we only cover three simple categories of reasoning relations: arithmetic, temporal, and world knowledge, each with shallow composition. While the coverage of reasoning types is limited, the core finding remains: current retrievers struggle to locate relevant documents when reasoning is implicit, and LLMs fail to reliably attend to the correct evidence in long-context settings.
Appendix A Dataset Generation
In this appendix, we describe, for each of the three reasoning categories, Arithmetic, World-Knowledge, and Temporal, (i) how we construct the implicit-tuple sets and (ii) what arguments are required to synthesize their corresponding contexts. After covering tuple-set construction, we explain in Section A.4 how a language model is used to generate the final passages.
A.1 Arithmetic Reasoning
For each implicit tuple set in the collection, to generate tuples for the Arithmetic category, we use Algorithm 1, which ensures that all tuples satisfy the these constraints: (i) all and values across the tuples are distinct, ensuring exactly one correct answer per query; (ii) all and values should be a multiple of 10, so that values resemble realistic prices; (iii) the values are evenly distributed across a predefined range of plausible prices, avoiding value clustering; and (iv) the resulting multiplier must have at most 2 decimals; hence, the required calculation is simple. The Algorithm1 returns a set of tuples of the form , where the two prices are mutually distinct and the multiplier satisfies predefined numerical constraints. The construction guarantees uniform distribution across price ranges and ensures that each tuple encodes a plausible relative-price comparison suitable for reasoning-based retrieval.
A.2 World-Knowledge Reasoning
As described in Section 2, we gather landmark-country pairs under three constraints: (i) each landmark must refer to a globally unique location, avoiding names that could correspond to multiple places; (ii) the landmark name must not include lexical, semantic, or language-specific cues that reveal its country, avoiding surface-form shortcuts; and (iii) landmarks must refer to specific, recognizable sites rather than generic ones. To do so, we first assemble a seed list of unique landmark-country pairs via five steps: (i) issue a SPARQL query to Wikidata to retrieve every entity whose instance-of (P31) chain includes exactly one of the high-level place classes, museum (Q33506), university (Q3918), church building (Q16970), venue (Q17350442), or landmark (the superclass of Q17350442), and that is linked to exactly one sovereign state via the country (P17) property; (ii) for each hit, extract the English labels of its enclosing administrative region (P131), city (P131 restricted to Q515), generic location (P276), and street (P669), yielding up to five concentric location strings; (iii) discard entities with missing, machine-generated, or multi-country labels, then drop any whose name tokens overlap these location strings; embed each remaining landmark-country pair with the 768-dimensional Contriever encoder and retain only those with cosine similarity below 0.25; (iv) pass the remainders to a 70B-parameter Llama-3.3 classifier that flags and removes generic names (e.g., “Downtown Club”) or labels leaking their country, using exponential back-off retries until accepted; and (v) submit each accepted label to GPT-4o in web-search mode666Checkpoint: gpt-4o-search-preview-2025-03-11, prompting it to return the place’s country in a dictionary format, and keep only those for which the model returns exactly one location. This process yields a balanced pool of 100 unique landmark-country pairs, ensuring we sample across different countries rather than multiple landmarks from the same country. Finally, for each implicit tuple set, we select a set of distinct countries and, for each , sample one landmark from that country’s filtered list . The resulting implicit tuple sets is as follows:
Generating implicit tuple sets, we have our collection.
A.3 Temporal Reasoning
As described in Section 2, for each tuple set, we generate a implicit tuple sets of tuples to have the following constraints: (i) all explicit dates in any are unique within the tuple set, ensuring that each query maps to exactly one positive document; (ii) all dates used as base or resolved targets are evenly distributed across a fixed date window to avoid clustering; and (iii) all relative offsets in must fall within a limited number of days from. Because context synthesis differs between multi-speaker and uni-speaker modes, we describe the two procedures separately in the following subsections.
A.3.1 Multi-Speaker: Tuple Construction
Here, we generate implicit tuple lists, each containing tuples. For generating them, we use Algorithm 2. The returned output contains all the information detailed before.
A.3.2 Uni-Speaker: Tuple Construction
We want to generate implicit tuple sets, each containing tuples. Each tuple set describes a main conversant’s 28-day activity schedule. We categorize activities into three types:
1.
One-time: executed exactly once in the 14-day period (we have 9 for then in each schedule).
2.
Repeating-Non-Sequential: occurring on multiple, non-consecutive days (we have 3 for them in each schedule: 2 days, 3 days, and 2 days).
3.
Repeating-Sequential: performed on consecutive days (we have 3 for them in each schedule: 3 days, 3 days, and 4 days).
Each set covers two consecutive 14-day blocks, contains exactly =30 activities, and is constructed in three phases: (i) scheduling activities without temporal overlap, (ii) selecting one message time per activity such that it differs from the scheduled lot, and (iii) packaging every activity into a tuple . Algorithm 3 details the procedure.
A.4 Synthesizing the Context
Now, for each reasoning category and document style, we have a list of implicit tuple sets containing all the information needed to have a consistent dataset. Consider having the auxiliary lexical content (personal names, daily-work verbs, brand names with corresponding items, and per-category forum topics and questions needed)777Generated using GPT-o1 OpenAI (2024), to each tuple of the implicit tuple list, we assign unique entities and then each tuple, contains all the information to generating the document in natural way. The exact item required for each reasoning category and style is mentioned in Table 4. Depending on the style, we proceed as follows:
A.4.1 Uni-Speaker (Chat-Style)
In the conversation-generation stage, we load the implicit tuple sets for each (category, style) pair. First, we use the STARTING_CONVERSATION_PROMPT to generate unique “starting phrases” for the tuple set (one for each tuple), ensuring that no two dialogues begin identically (Figure 2). Next, for each tuple, we prepend one of these starting phrases and feed the combination into the CONVERSATION_GENERATION_PROMPT; category-specific prompt templates and requirements are shown in Figure 5 for the Arithmetic, Figure 9 for the World-Knowledge, and Figure 13 for the Temporal reasoning. We then ask the model to produce exactly ten utterances per chat, verify the count, and regenerate any that do not meet this criterion. Finally, each completed conversation is submitted to a separate LLM via the FEATURE_EXTRACTION_PROMPT, which must reconstruct the original tuple to confirm that the dialogue faithfully conveys the intended information (Figure 7 for the Arithmetic, Figure 11 for the World-Knowledge, and Figure 15 for the Temporal reasoning).
A.4.2 Multi-Speaker (Forum-Style)
In the forum-style generation stage, we similarly load the implicit tuple sets for each (category, style) pair. We first generate unique starting phrases for the tuple set using the STARTING_CONVERSATION_PROMPT, so that each reply begins differently (Figure 4 for the Arithmetic, Figure 8 for the World-Knowledge, and Figure 12 for the Temporal reasoning). Then, for each tuple, we provide the forum topic, its base question, one starting phrases, and the tuple data to the CONVERSATION_GENERATION_PROMPT configured for forum responses; category-specific templates and requirements again appear in Figure 5 for the Arithmetic, Figure 8 for the World-Knowledge, and Figure 13 for the Temporal reasoning. We generate exactly five sentences per response. Finally, each forum reply is passed to a separate LLM via the FEATURE_EXTRACTION_PROMPT to extract the original tuple, ensuring the response accurately encodes the tuple’s information (Figure 6 for the Arithmetic, Figure 10 for the World-Knowledge, and Figure 14 for the Temporal reasoning).
Reasoning | Multi-Speaker | Uni-Speaker | ||
Each Pool | Each Document | Each Pool | Each Document | |
Arithmetic | Topic, | Conversant, Brand, Model Years (Two random numbers | Main Conversant | Second Conversant, Shopping |
Forum-Base-Question | between 2013 up to 2024, the lower number is | item, Low priced brand | ||
the price of the lower-priced item) | name, High priced brand | |||
World-Knowledge | Topic, | Conversant, | Main Conversant | Second Conversant, Landmark, |
Forum-Base-Question | Landmark | Message Date | ||
Temporal | Topic, | Conversant, Item related to the Forum Base Question, | Main Conversant | Second Conversant, |
Forum-Base-Question | Daily work verb |
Appendix B Results
Retriever | Reasoning | Average | ||
W. Know. | Arithmetic | Temporal | ||
Sparse Baseline | ||||
BM25 | 9.15 | 7.42 | 7.20 | 7.92 |
Late-Interaction | ||||
ColBERT v2 | 10.75 | 9.61 | 7.90 | 9.42 |
Dense Encoders | ||||
Contriever | 10.85 | 8.97 | 8.18 | 9.33 |
Dragon+ | 11.80 | 9.53 | 8.30 | 9.88 |
ReasonIR | 14.67 | 7.04 | 10.14 | 10.62 |
Knowledge-Graph–Augmented Indexer | ||||
HippoRAG 2 | 11.11 | 9.36 | 8.13 | 9.54 |
As explained in Section 4, our main retrieval metric is nDCG@10. For completeness, we also compute MRR@10, summarized in Table 5. The lower MRR@10 scores confirm that, across reasoning categories, the systems often fail to rank the positive documents first, underscoring the modest overall performance already suggested by nDCG. Granular results for the Uni-Speaker and Multi-Speaker settings are provided in Table 6. For the RAG-style evaluation, model outputs were generated using the prompt templates shown in Figures 16 and 17, and then evaluated using ROUGE-1 Recall against the reference answer.
Experiment | World-Knowledge | Arithmetic | Temporal | Average | ||||||||||||
Uni-Speaker | Multi-Speaker | Uni-Speaker | Multi-Speaker | Uni-Speaker | Multi-Speaker | Uni-Speaker | Multi-Speaker | |||||||||
MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | |
Sparse Baseline | ||||||||||||||||
BM25 | 8.44 | 12.98 | 9.85 | 15.22 | 9.23 | 14.05 | 5.61 | 8.07 | 7.06 | 10.68 | 7.33 | 11.76 | 8.24 | 12.57 | 7.60 | 11.68 |
Late-Interaction | ||||||||||||||||
ColBERT v2 | 10.07 | 15.12 | 11.43 | 16.95 | 9.49 | 14.63 | 9.73 | 15.24 | 7.54 | 11.29 | 8.26 | 13.11 | 9.03 | 13.68 | 9.81 | 15.10 |
Dense Encoders | ||||||||||||||||
Contriever | 9.88 | 14.80 | 11.81 | 17.51 | 8.49 | 13.17 | 9.45 | 14.51 | 7.58 | 11.91 | 8.79 | 13.27 | 8.65 | 13.29 | 10.01 | 15.09 |
Dragon+ | 10.00 | 14.92 | 13.59 | 19.38 | 9.24 | 14.10 | 9.83 | 15.11 | 7.30 | 10.93 | 9.31 | 14.13 | 8.85 | 13.32 | 10.91 | 16.21 |
ReasonIR | 8.12 | 11.50 | 21.21 | 27.57 | 3.71 | 5.66 | 10.37 | 15.81 | 7.70 | 11.51 | 12.59 | 18.35 | 6.51 | 9.56 | 14.72 | 20.58 |
Knowledge-Graph–Augmented Indexer | ||||||||||||||||
HippoRAG 2 | 10.26 | 15.10 | 11.97 | 17.66 | 8.47 | 13.17 | 10.26 | 15.26 | 7.27 | 11.54 | 9.00 | 13.45 | 8.66 | 13.27 | 10.41 | 15.46 |

### v2

@@
 In contrast, we present ImpliRet, a benchmark that shifts the reasoning challenge to document-side processing: The queries are simple, but relevance depends on facts stated implicitly in documents through temporal (e.g., resolving “two days ago”), arithmetic, and world knowledge relationships.
-We evaluate a range of sparse and dense retrievers, all of which struggle in this setting: the best nDCG@10 is only 15.07%.
+We evaluate a range of sparse and dense retrievers, all of which struggle in this setting: the best nDCG@10 is only 14.91%.
 We also test whether long-context models can overcome this limitation.
-But even with a short context of only ten documents, including the positive document, GPT-4.1 scores only 35.06%, showing that document-side reasoning remains a challenge.
+But even with a short context of only thirty documents, including the positive document, GPT-o4-mini scores only 55.54%, showing that document-side reasoning remains a challenge.
 Our codes are available at github.com/ZeinabTaghavi/IMPLIRET.††* Equal Contribution.
@@
 (2022), and Dragon+ Lin et al.
-(2023), and observe consistently poor performance: the best nDCG@10 is only 15.07% across our benchmark.
+(2023), and observe consistently poor performance: the best nDCG@10 is only 14.91% across our benchmark.
 To test whether long-context capabilities could mitigate the problem, we evaluate models in a setting where the positive document is included among several distractors.
-While GPT-4.1 answers correctly when given only the positive document, its performance drops sharply even with just ten documents in-context, achieving a ROUGE-1 recall of 35.06%.
+While GPT-o4-mini answers correctly when given only the positive document, its performance drops sharply even with just thirty documents in-context, achieving a ROUGE-1 recall of 55.54%.
 Our dataset ImpliRet introduces a new setting that requires document-side reasoning for retrieval rather than query-side reasoning.
@@
 Tuples are generated using a sampling algorithm that selects base prices and checks constraint satisfaction, backtracking as needed until valid tuples are found (where is the target number of documents indicated as “Docs” in Table 1).
-Full constraint details and sampling logic are provided in Appendix 2.1.
+Full constraint details and sampling logic are provided in Appendix A.1.
 World Knowledge.
@@
 This pipeline is supported by auxiliary lexical resources, including random names, brand-item pairs, and activity lists, as well as per-reasoning category prompt templates.
-We use LLAMA 3.3-70B Meta (2024) to synthesize the documents for each tuple.111Details such as prompts and query templates are available in Appendix A.4.
+We use Gemma-3-27B-it Team (2025) to synthesize the documents for each tuple.111Details such as prompts and query templates are available in Appendix A.4.
 Table 1 presents ImpliRet statistics222The tokens are counted using GPT-2 tokenizer Radford et al.
@@
 Sparse | ||||
-BM25 | 14.10 | 11.06 | 11.22 | 12.13 |
+BM25 | 14.69 | 11.06 | 10.98 | 12.24 |
 Late-Interaction | ||||
-ColBERT v2 | 16.04 | 14.93 | 12.20 | 14.39 |
+ColBERT v2 | 15.79 | 14.96 | 11.99 | 14.25 |
 Dense Encoders | ||||
-Contriever | 16.15 | 13.84 | 12.59 | 14.19 |
-Dragon+ | 17.15 | 14.61 | 12.53 | 14.76 |
-ReasonIR | 19.53 | 10.74 | 14.93 | 15.07 |
+Contriever | 16.50 | 13.70 | 12.73 | 14.31 |
+Dragon+ | 17.46 | 14.61 | 12.66 | 14.91 |
+ReasonIR | 18.88 | 10.78 | 11.25 | 13.64 |
 Knowledge-Graph Augmented Indexer | ||||
-HippoRAG 2 | 16.38 | 14.21 | 12.50 | 14.36 |
+HippoRAG 2 | 16.62 | 14.13 | 12.83 | 14.53 |
 4 Results
 The nDCG@10 results across all reasoning categories are presented in Table 2.
-The highest average score, 15.07 (achieved by ReasonIR, a recent 8B-parameter LLM), shows the difficulty retrieval models face when reasoning over implicit facts in documents.
-More efficient baselines such as Contriever, Dragon+, and BM25 perform substantially worse; notably, BM25 reaches just 12.13 due to its reliance on surface-level lexical overlap.
-Performance varies across reasoning types: the World Knowledge category exhibits the largest performance spread (14.10 vs.
-19.53), while it is narrowest for Arithmetic (10.74 vs.
-14.61).
-Discourse style also plays a role: ReasonIR scores 20.58 on multi-speaker examples compared to 9.56 on uni-speaker ones, suggesting that stylistic structure affects retrieval difficulty.333Full results per category and style in Appendix B.
+The highest average score, 14.91 (achieved by Dragon+), shows the difficulty retrieval models face when reasoning over implicit facts in documents.
+More efficient baselines such as Contriever and BM25 perform substantially worse; notably, BM25 reaches just 12.24 due to its reliance on surface-level lexical overlap.
+Performance varies across reasoning types: the Arithmetic category exhibits the largest performance spread (14.96 vs.
+10.78), while it is narrowest for Temporal (12.83 vs.
+10.98).
+Discourse style also plays a role: Dragon+ scores 16.45% on multi-speaker examples compared to 13.37 on uni-speaker ones, suggesting that stylistic structure affects retrieval difficulty.333Full results per category and style in Appendix B.
 Experiment | Reasoning | Average | |||
@@
 10 | 27.37 | 16.98 | 25.23 | 23.19 | |
-All | 17.43 | 4.42 | 10.29 | 10.71 | |
+30 | 17.43 | 4.42 | 10.29 | 10.71 | |
 GPT-4.1 | 1 | 93.24 | 92.12 | 84.90 | 88.05 |
 10 | 62.21 | 23.86 | 15.59 | 35.06 | |
-All | 53.91 | 9.28 | 6.93 | 22.90 |
+30 | 53.91 | 9.28 | 6.93 | 22.90 | |
+GPT-o4-mini | 1 | 92.34 | 92.45 | 93.44 | 92.74 |
+10 | 88.11 | 76.61 | 73.94 | 79.55 | |
+30 | 75.44 | 76.31 | 14.86 | 55.54 |
 RAG Performance with an Oracle Retriever on Reason-Sensitive Documents.
@@
 This configuration removes retrieval as a variable and isolates the LLM’s document-side reasoning ability.
-We evaluate three settings: (positive only), (positive plus nine negatives), and a full-pool setting where all documents from the pool are provided as context.
+We evaluate three settings: (positive only), (positive plus nine negatives), and (a full-pool setting where all documents from the pool are provided as context).
 The model receives the query along with the sequence of documents and must generate an answer.
-We evaluate two reader models: LLAMA 3.3 70B and GPT-4.1.444Checkpoint: gpt-4.1-2025-04-14 In Table 3, we report the average ROUGE-1 recall555R-1 Rec.
+We evaluate three reader models: LLAMA 3.3 70B, GPT-4.1444Checkpoint: gpt-4.1-2025-04-14, and GPT-o4-mini555Checkpoint: o4-mini-2025-04-16.
+In Table 3, we report the average ROUGE-1 recall666R-1 Rec.
 scores to measure the overlap between the generated output and the positive answer Lin (2004).
-When given only the positive document (), the two models achieve average ROUGE-1 Recall of 81.92 and 88.05.
+When given only the positive document (), the models achieve average ROUGE-1 Recall of 81.92, 88.05, and 92.74.
 This suggests that the query itself is straightforward to answer once the relevant document is isolated.
@@
 It covers three reasoning types – world knowledge, arithmetic, and temporal – and two discourse styles.
-Across sparse, dense, and KG-augmented retrievers, the best nDCG@10 is only 15.07.
-Even with GPT-4.1, given ten documents including the gold, performance peaks at just 35.06%.
+Across sparse, dense, and KG-augmented retrievers, the best nDCG@10 is only 14.91.
+Even with GPT-o4-mini, given thirty documents including the gold, performance peaks at just 55.54%.
 These results highlight the difficulty of retrieving implicit facts and the need for models that can reason beyond surface cues.
@@
 As described in Section 2, we gather landmark-country pairs under three constraints: (i) each landmark must refer to a globally unique location, avoiding names that could correspond to multiple places; (ii) the landmark name must not include lexical, semantic, or language-specific cues that reveal its country, avoiding surface-form shortcuts; and (iii) landmarks must refer to specific, recognizable sites rather than generic ones.
-To do so, we first assemble a seed list of unique landmark-country pairs via five steps: (i) issue a SPARQL query to Wikidata to retrieve every entity whose instance-of (P31) chain includes exactly one of the high-level place classes, museum (Q33506), university (Q3918), church building (Q16970), venue (Q17350442), or landmark (the superclass of Q17350442), and that is linked to exactly one sovereign state via the country (P17) property; (ii) for each hit, extract the English labels of its enclosing administrative region (P131), city (P131 restricted to Q515), generic location (P276), and street (P669), yielding up to five concentric location strings; (iii) discard entities with missing, machine-generated, or multi-country labels, then drop any whose name tokens overlap these location strings; embed each remaining landmark-country pair with the 768-dimensional Contriever encoder and retain only those with cosine similarity below 0.25; (iv) pass the remainders to a 70B-parameter Llama-3.3 classifier that flags and removes generic names (e.g., “Downtown Club”) or labels leaking their country, using exponential back-off retries until accepted; and (v) submit each accepted label to GPT-4o in web-search mode666Checkpoint: gpt-4o-search-preview-2025-03-11, prompting it to return the place’s country in a dictionary format, and keep only those for which the model returns exactly one location.
+To do so, we first assemble a seed list of unique landmark-country pairs via five steps: (i) issue a SPARQL query to Wikidata to retrieve every entity whose instance-of (P31) chain includes exactly one of the high-level place classes, museum (Q33506), university (Q3918), church building (Q16970), venue (Q17350442), or landmark (the superclass of Q17350442), and that is linked to exactly one sovereign state via the country (P17) property; (ii) for each hit, extract the English labels of its enclosing administrative region (P131), city (P131 restricted to Q515), generic location (P276), and street (P669), yielding up to five concentric location strings; (iii) discard entities with missing, machine-generated, or multi-country labels, then drop any whose name tokens overlap these location strings; embed each remaining landmark-country pair with the 768-dimensional Contriever encoder and retain only those with cosine similarity below 0.25; (iv) pass the remainders to a 70B-parameter Llama-3.3 classifier that flags and removes generic names (e.g., “Downtown Club”) or labels leaking their country, using exponential back-off retries until accepted; and (v) submit each accepted label to GPT-4o in web-search mode777Checkpoint: gpt-4o-search-preview-2025-03-11, prompting it to return the place’s country in a dictionary format, and keep only those for which the model returns exactly one location.
 This process yields a balanced pool of 100 unique landmark-country pairs, ensuring we sample across different countries rather than multiple landmarks from the same country.
@@
 Now, for each reasoning category and document style, we have a list of implicit tuple sets containing all the information needed to have a consistent dataset.
-Consider having the auxiliary lexical content (personal names, daily-work verbs, brand names with corresponding items, and per-category forum topics and questions needed)777Generated using GPT-o1 OpenAI (2024), to each tuple of the implicit tuple list, we assign unique entities and then each tuple, contains all the information to generating the document in natural way.
+Consider having the auxiliary lexical content (personal names, daily-work verbs, brand names with corresponding items, and per-category forum topics and questions needed)888Generated using Gemma-3-27B-it Team (2025), to each tuple of the implicit tuple list, we assign unique entities and then each tuple, contains all the information to generating the document in natural way.
 The exact item required for each reasoning category and style is mentioned in Table 4.
@@
 Sparse Baseline | ||||
-BM25 | 9.15 | 7.42 | 7.20 | 7.92 |
+BM25 | 9.55 | 7.42 | 6.83 | 7.93 |
 Late-Interaction | ||||
-ColBERT v2 | 10.75 | 9.61 | 7.90 | 9.42 |
+ColBERT v2 | 10.59 | 9.63 | 7.55 | 9.26 |
 Dense Encoders | ||||
-Contriever | 10.85 | 8.97 | 8.18 | 9.33 |
-Dragon+ | 11.80 | 9.53 | 8.30 | 9.88 |
-ReasonIR | 14.67 | 7.04 | 10.14 | 10.62 |
+Contriever | 11.19 | 8.84 | 8.48 | 9.50 |
+Dragon+ | 11.97 | 9.47 | 8.26 | 9.90 |
+ReasonIR | 14.23 | 7.13 | 7.78 | 9.71 |
 Knowledge-Graph–Augmented Indexer | ||||
-HippoRAG 2 | 11.11 | 9.36 | 8.13 | 9.54 |
+HippoRAG 2 | 11.30 | 9.28 | 8.57 | 9.72 |
 As explained in Section 4, our main retrieval metric is nDCG@10.
@@
 Uni-Speaker | Multi-Speaker | Uni-Speaker | Multi-Speaker | Uni-Speaker | Multi-Speaker | Uni-Speaker | Multi-Speaker | |||||||||
-MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | |
+MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | |
 Sparse Baseline | ||||||||||||||||
-BM25 | 8.44 | 12.98 | 9.85 | 15.22 | 9.23 | 14.05 | 5.61 | 8.07 | 7.06 | 10.68 | 7.33 | 11.76 | 8.24 | 12.57 | 7.60 | 11.68 |
+BM25 | 9.23 | 14.18 | 9.86 | 15.20 | 9.23 | 14.05 | 5.61 | 8.07 | 6.33 | 10.20 | 7.33 | 11.76 | 8.26 | 12.81 | 7.60 | 11.68 |
 Late-Interaction | ||||||||||||||||
-ColBERT v2 | 10.07 | 15.12 | 11.43 | 16.95 | 9.49 | 14.63 | 9.73 | 15.24 | 7.54 | 11.29 | 8.26 | 13.11 | 9.03 | 13.68 | 9.81 | 15.10 |
+ColBERT v2 | 9.57 | 14.54 | 11.60 | 17.03 | 9.45 | 14.62 | 9.81 | 15.30 | 6.85 | 10.86 | 8.25 | 13.12 | 8.62 | 13.34 | 9.89 | 15.15 |
 Dense Encoders | ||||||||||||||||
-Contriever | 9.88 | 14.80 | 11.81 | 17.51 | 8.49 | 13.17 | 9.45 | 14.51 | 7.58 | 11.91 | 8.79 | 13.27 | 8.65 | 13.29 | 10.01 | 15.09 |
-Dragon+ | 10.00 | 14.92 | 13.59 | 19.38 | 9.24 | 14.10 | 9.83 | 15.11 | 7.30 | 10.93 | 9.31 | 14.13 | 8.85 | 13.32 | 10.91 | 16.21 |
-ReasonIR | 8.12 | 11.50 | 21.21 | 27.57 | 3.71 | 5.66 | 10.37 | 15.81 | 7.70 | 11.51 | 12.59 | 18.35 | 6.51 | 9.56 | 14.72 | 20.58 |
+Contriever | 9.77 | 14.77 | 12.60 | 18.23 | 8.23 | 12.87 | 9.45 | 14.52 | 8.17 | 12.22 | 8.78 | 13.25 | 8.72 | 13.29 | 10.28 | 15.33 |
+Dragon+ | 9.94 | 14.85 | 14.00 | 20.07 | 9.11 | 14.11 | 9.83 | 15.12 | 7.20 | 11.16 | 9.32 | 14.16 | 8.75 | 13.37 | 11.05 | 16.45 |
+ReasonIR | 7.75 | 10.54 | 20.71 | 27.21 | 3.87 | 5.68 | 10.39 | 15.89 | 2.95 | 4.17 | 12.60 | 18.33 | 4.86 | 6.80 | 14.57 | 20.48 |
 Knowledge-Graph–Augmented Indexer | ||||||||||||||||
-HippoRAG 2 | 10.26 | 15.10 | 11.97 | 17.66 | 8.47 | 13.17 | 10.26 | 15.26 | 7.27 | 11.54 | 9.00 | 13.45 | 8.66 | 13.27 | 10.41 | 15.46 |
+HippoRAG 2 | 9.95 | 14.94 | 12.66 | 18.30 | 8.31 | 12.97 | 10.26 | 15.29 | 8.18 | 12.24 | 8.96 | 13.42 | 8.81 | 13.38 | 10.63 | 15.67 |

### v3

@@
 ImpliRet: Benchmarking the Implicit Fact Retrieval Challenge
-Zeinab Sadat Taghavi*, Ali Modarressi*, Yunpu Ma, Hinrich Schütze Center for Information and Language Processing, LMU Munich Munich Center for Machine Learning (MCML) Correspondence: zeinabtaghavi@cis.lmu.de
+Zeinab Sadat Taghavi*, Ali Modarressi*, Yunpu Ma, Hinrich Schütze Center for Information and Language Processing, LMU Munich Munich Center for Machine Learning (MCML) Correspondence: [zeinabtaghavi, amodaresi]@cis.lmu.de
 1 Introduction
@@
 | Temporal | Uni-speaker | 1500 | 479 | 719,226 |
-Multi-speaker | 1500 | 141 | 212,502 |
+| Multi-speaker | 1500 | 141 | 212,502 | |
 2.2 Document-Query Pairs
@@
 (2019)..
+Fluency and implicitness sanity check.
+We drew a stratified random sample of 72 instances (query–document pairs; 3 reasoning categories × 2 discourse styles × 12 per cell) and manually assessed each for (i) fluency, (ii) implicit support for the queried fact, and (iii) absence of explicit leakage (i.e., a verbatim statement of the fact).
+In this sample, all documents were fluent and supported the queries implicitly; under our rubric, we observed no cases of explicit leakage.
+Further details are provided in Appendix A.5.
 3 Experiments
@@
 | Llama 3.3 70B | 1 | 73.79 | 90.13 | 81.85 | 81.92 |
-10 | 27.37 | 16.98 | 25.23 | 23.19 | |
-30 | 17.43 | 4.42 | 10.29 | 10.71 | |
+| 10 | 27.37 | 16.98 | 25.23 | 23.19 | |
+| 30 | 17.43 | 4.42 | 10.29 | 10.71 | |
 | GPT-4.1 | 1 | 93.24 | 92.12 | 84.90 | 88.05 |
-10 | 62.21 | 23.86 | 15.59 | 35.06 | |
-30 | 53.91 | 9.28 | 6.93 | 22.90 | |
+| 10 | 62.21 | 23.86 | 15.59 | 35.06 | |
+| 30 | 53.91 | 9.28 | 6.93 | 22.90 | |
 | GPT-o4-mini | 1 | 92.34 | 92.45 | 93.44 | 92.74 |
-10 | 88.11 | 76.61 | 73.94 | 79.55 | |
-30 | 75.44 | 76.31 | 14.86 | 55.54 |
+| 10 | 88.11 | 76.61 | 73.94 | 79.55 | |
+| 30 | 75.44 | 76.31 | 14.86 | 55.54 | |
 RAG Performance with an Oracle Retriever on Reason-Sensitive Documents.
@@
 (2025).
+Error Analysis
+RAG has two stages—retrieval and generation—so we analyze errors along two axes.
+1.
+Retrieval side (Rank-1 vs.
+Positive).
+For each query, we compare the retriever’s top-1 passage with the annotated positive.
+We analyze the 60 queries where Dragon+’s top-1 document differs from the positive ( reasoning categories discourse styles queries), yielding 120 passages (top-1 and positive per query).
+We categorize mis-rank reasons into four groups: (i) Word Overlap (top-1 has extra query surface tokens), (ii) Semantic Cue (similar overlap but extra topical/theme terms), (iii) Length (overlap/semantics comparable; shorter passage chosen), and (iv) Unknown (indistinguishable under our heuristics).
+Table 4 shows that Semantic Cue is most frequent in arithmetic queries, and Word Overlap in temporal and world knowledge.
+2.
+Generation side (Oracle-RAG, vs.
+all).
+To isolate generation errors, we evaluate an oracle setting where the positive passage is guaranteed in context.
+We randomly select 60 queries () and evaluate two context sizes ( and , others selected randomly), yielding 120 cases.
+After reviewing outputs, we assign a single label to each incorrect answer: (i) Malformation (positive present, answer malformed), (ii) No-answer/unrelated, or (iii) Distraction (copied from a surface-similar distractor).
+Table 5 shows that No-answer/unrelated is most frequent, Distraction occurs mainly in temporal queries, and longer context reduces Malformation but raises No-answer/unrelated, suggesting that context length alone does not fix generation errors.
+| Reasoning | Word overlap | Semantic cue | Length | Unknown |
+| Arithmetic | 15% | 55% | 5% | 25% |
+| Temporal | 55% | 5% | 35% | 5% |
+| W. Know. | 50% | 5% | 30% | 15% |
+| Reasoning | Malformation | No-answer/Unrelated | Distraction | |
+| Arithmetic | 10 | 40% | 60% | 0% |
+| all | 20% | 75% | 5% | |
+| W. Know. | 10 | 45% | 55% | 0% |
+| all | 15% | 85% | 0% | |
+| Temporal | 10 | 0% | 35% | 65% |
+| all | 0% | 40% | 60% |
 5 Conclusion
@@
 Across sparse, dense, and KG-augmented retrievers, the best nDCG@10 is only 14.91.
-Even with GPT-o4-mini, given thirty documents including the gold, performance peaks at just 55.54%.
+Even with GPT-o4-mini, given thirty documents including the positive, performance peaks at just 55.54%.
 These results highlight the difficulty of retrieving implicit facts and the need for models that can reason beyond surface cues.
@@
 Finally, each forum reply is passed to a separate LLM via the FEATURE_EXTRACTION_PROMPT to extract the original tuple, ensuring the response accurately encodes the tuple’s information (Figure 6 for the Arithmetic, Figure 10 for the World-Knowledge, and Figure 14 for the Temporal reasoning).
+A.5 Human Evaluation Details
+We performed a small, stratified sanity check to complement automatic validation.
+For each reasoning category discourse style cell, we randomly selected two queries (12 instances per cell).
+For each selected query, candidate passages were ranked with the ReasonIR retriever, and six passages were drawn by uniform sampling from two rank strata: top–20 (3) and bottom–20 (3).
+This yielded 72 passages in total.
+We assessed each passage for (i) fluency, (ii) implicit support of their corresponding queried fact (entailed but not stated verbatim), and (iii) absence of explicit leakage.
+In this sample, all passages were fluent and passed the implicitness and non-leakage checks under our rubric.
 | Reasoning | Multi-speaker | Uni-speaker | ||
@@
 | Temporal | Topic, | Conversant, Item related to the Forum Base Question, | Main Conversant | Second Conversant, |
-Forum-Base-Question | Daily work verb |
+| Forum-Base-Question | Daily work verb | |||
 Appendix B Results